Today (T=0), an investor purchased a nine year bond with an 8.0% coupon for $9,680. The...

  1. Today (T=0), an investor purchased a nine year bond with an 8.0% coupon for $9,680. The bond has a face value of $10,000. In six months (T=0.5) interest rates have increased by 1.0% and the investor decides to sell the bond immediately after receiving the first coupon payment. What is the investor’s total gain (loss) on the bond? HINT: Total Gain (Loss) = Price Change in Bond + Coupon

    A. ($583) B. ($183) C. ($150) D. $190 E. $990
Answer #1

Coupon = 8%*10000 = 800
Coupon semiannual = 800/2 = 400
For YTM using Financial Calculator we get
N = 18; PMT = 400; FV = 10000; PV = -9680; CPT I/Y
YTM half Yearly = 4.246%

Rate = 4.246%*2 = 8.492%

New YTM = 8.52%+1 = 9.492%
New Price of bond = PV of Coupons for 17 months + PV of Par Value = 400*(1-(1+9.492%)-17/9.492% + 10000/(1+9.492%)17 = 9127.60

Loss= 9126.60 + 400 - 9680 = -151(approx)